May 26, 2011

Data saved in /data/luci/20110527

Observers: Marie Rex, Emily Mclinden, Marco Pedani

TO: Geno

Scripts observed saved on lucifer in: /home/observer/AZ/May2011/

Data saved in /data/luci/20110528

7:30pm tentatively open --high winds

7:45pm starting Emily's first target --into the wind (~15m/s)

8:00pm winds are high, took too long to collimate, resending preset

8:10pm telescope control computer was not registering Emily's preset, had to cancel and resend 3 times

Seeing ~1.2"

8:23pm starting Emily's science scripts

8:50pm starting Emily's calibration arcs/flats

Had to move calibration unit by hand

9:18pm going to Emily's second target --34 deg, too low

9:26pm going to one of her higher elevation targets

Seeing 0.6--0.8

10:23pm doing flats, same calibration unit problem

10:30pm going to Emily's third target

Seeing 1" --2"

11:00pm calibration flats

11:10pm Acquisition for Rex target A1914

#98 -- 101

11:25pm Starting H-band scirnce script for A1914

#102-- 125 (Seeing 0.8" -- 1.5", frame 103 ~2")

1:40am Doing flats

#126 --#135

1:50am telluric acquisition for A1914

#136--139

2:00am H-band telluric measurement

#140 --143

2:05am K-band telluric measurement

#144 --147

2:10am Going back to A1914

#148 --151

2:20am K-band science for A1914

#152 --163 (seeing 1--2")

3:30am K-band flats

#164 --173

4:00am stating imaging script for MACS 2043

#174--281

4:46am Running EndOfNight

5:00 Running DoDarks.txt

-- XiaohuiFan - 27 May 2011
Topic revision: r5 - 28 May 2011, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback